What current will deposit 3.25grams of zinc for 2 hours


1
Expert's answer
2021-08-06T05:15:27-0400

Solution:

Faraday's law of electrolysis can be used.

Faraday's law of electrolysis can be expressed as:

m = (Q × M) / (Z × F) = (I × t × M) / (Z × F)

where:

m = mass of material (g)

M = molar mass of material (g/mol)

I = current (A)

t = time (s)

Z = number of electrons

F = Faraday constant, 96485 (C mol-1)


Therefore,

I = (m × Z × F) / (t × M)

where:

m(Zn) = 3.25 g

Z = 2 (Zn2+ + 2e → Zn)

F = 96485 C mol-1

t = (2 h) × (60 min / 1 h) × (60 s / 1 min) = 7200 s

M(Zn) = 65 g/mol

Plug in the numbers and solve for I:

I = (3.25 g × 2 × 96485 C mol-1) / (7200 s × 65 g/mol) = 1.34 C/s = 1.34 A

I = 1.34 A


Answer: 1.34 A